OpenStudy (anonymous):

Find an exact simplified solution to the equation on the interval 0 < x < 1. 1 = 6cos(x + 1) − 2

OpenStudy (dumbcow):

just try to solve for cos what do you get ? cos(x+1) = ??

OpenStudy (anonymous):

i got basically this 1/2= cos(x+1) so cos = pi/3 or 5pi/3 but it is not in the set parameter

OpenStudy (dumbcow):

very good this is because you forgot to subtract 1 :) x+1 = pi/3 x = pi/3 - 1

OpenStudy (dumbcow):

5pi/3 - 1 is outside parameters of course

OpenStudy (anonymous):

oh ok thank you so much

OpenStudy (dumbcow):

yw just remember when you get the angle you still have to set it equal to whatever is in parenthesis